OpenStudy (anonymous):

Which equation is a quadratic equation? 3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3) I think it's y = 7x(5 − 2x) + 9 because it looks like it's in standard form but I just want to make sure.

OpenStudy (nubeer):

yes..

OpenStudy (johnweldon1993):

You would be correct :)

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

hre are more than one quadratic equations above.

OpenStudy (nubeer):

nop.. there are not.

OpenStudy (johnweldon1993):

@sami-21 no...that last equation would be a cubic

OpenStudy (anonymous):

7x2 + y = (3 + x)(7x − 1) y = 7x(5 − 2x) + 9

OpenStudy (nubeer):

7x^2 cancels up if u evaluate this option

OpenStudy (anonymous):

ok its one

OpenStudy (johnweldon1993):

correct @nubeer